OpenStudy (anonymous):

anyone (6-sqrt(3) ) (6+sqrt(3) =

OpenStudy (anonymous):

\[(6-\sqrt{3})(6+\sqrt{3})=36-3=33\]

OpenStudy (anonymous):

thanks a bunch how about sqrt(3)/sqrt(12)-sqrt(3)